# Tel.: +91-080-2293 2058

### v1

@@ Abstract
Initial boundary value problems arise commonly in applications with engineering and natural systems governed by nonlinear partial differential equations (PDEs). Operator learning is an emerging field for solving these equations by using a neural network to learn a map between infinite dimensional input and output function spaces. These neural operators are trained using a combination of data (observations or simulations) and PDE-residuals (physics-loss). A major drawback of existing neural approaches is the requirement to retrain with new initial/boundary conditions, and the necessity for a large amount of simulation data for training. We develop a physics-informed transformer neural operator (named PINTO) that efficiently generalizes to unseen initial and boundary conditions, trained in a simulation-free setting using only physics loss. The main innovation lies in our new iterative kernel integral operator units, implemented using cross-attention, to transform the PDE solution’s domain points into an initial/boundary condition-aware representation vector, enabling efficient learning of the solution function for new scenarios. The PINTO architecture is applied to simulate the solutions of important equations used in engineering applications: advection, Burgers, and steady and unsteady Navier-Stokes equations (three flow scenarios). For these five test cases, we show that the relative errors during testing under challenging conditions of unseen initial/boundary conditions are only one-fifth to one-third of other leading physics informed operator learning methods. Moreover, our PINTO model is able to accurately solve the advection and Burgers equations at time steps that are not included in the training collocation points. The code is available at https://github.com/quest-lab-iisc/PINTO.
@@
Nonlinear mapping Operator learning Transformer operator Scientific machine learning Physics-informed deep learning Initial and boundary generalization AI for PDEs
journal: myjournal
@@ 1 Introduction
Solving initial boundary value problems (IBVPs) with governing partial differential equations (PDEs) is essential to understand and characterize the dynamics of natural and engineering systems. Traditionally, numerical methods (e.g., finite difference, finite element, finite volume, or spectral) are employed to solve PDEs. In recent decades, scientists have created high-performance computing software designed to solve IBVP and parametric PDE, enabling their use as forward models in engineering design (Wang et al., 2022a; Ion et al., 2022; Richter et al., 2021; Heermann and Burkitt, 2012). However, for a new set of initial and/or boundary conditions, the numerical method must be applied afresh to solve the PDE as a new problem by spending significant compute resources. In contrast, contemporary artificial intelligence tasks in computer vision and natural language processing involve running a trained model (inference), which is computationally less expensive, for new scenarios. Those fields have benefitted from neural models that generalize to new situations well. The promise and application scope of such PDE solving models that could be trained once and used for fast simulation in new situations is appealing. Thus, over the past several years, the development of neural models for simulating systems governed by PDEs has picked up pace. Our goal is to develop a neural model for learing generalized solutions of IBVPs.
Deep neural networks that solve PDEs are broadly classified into two categories: (i) neural operators that learn the mapping between function spaces (input function of say parameters and output function of the solution) such as DeepONets, Fourier neural operators (FNO), physics-informed neural operators (PINO), and graph neural operators (GNO) (Li et al., 2020b, a; Lu et al., 2021a; You et al., 2022a; Kovachki et al., 2023a) and (ii) neural networks that learn a map between the PDE’s space-time domain and the solution function (Raissi et al., 2019; Lu et al., 2021b; Karniadakis et al., 2021; Yang et al., 2021; Cai et al., 2021). Both approaches can be trained using a combination of simulation data and PDE-residuals. When training is dominated by PDE-residuals (also called physics loss), the approach is called unsupervised, and a qualifier of physics-informed is added (e.g., Physics-Informed Neural Operator: PINO; Physics-Informed Neural Network: PINN). The physics loss is calculated using automatic differentiation to compute the partial derivatives of the solution with respect to the coordinates of the space-time domain. However, it must be noted that supervised training with simulation data obtained from a classical numerical method is also sometimes called physics-informed.
@@
A PINN model only learns the map for one instance of the PDE and requires re-training for solving with new initial/boundary conditions, much like traditional numerical solvers. On the other hand, neural operators can learn a map between an initial/boundary function and a solution function, thereby achieving generalization to unseen conditions for a particular PDE. However, the current leading neural operators such as the Fourier neural operator (FNO) (Li et al., 2020a; Pathak et al., 2022; Wen et al., 2022; Bonev et al., 2023; You et al., 2022b; Li et al., 2023b; George et al., 2024; Kovachki et al., 2021a, 2023b; Lehmann et al., 2024), DeepONet (Wang et al., 2021; Lu et al., 2021a; Lanthaler et al., 2022; He et al., 2024; Xu et al., 2023; Li et al., 2023a; Goswami et al., 2022, 2023, 2024; Haghighat et al., 2024), Graph neural operator (Li et al., 2020b, c), Convolutional neural operators (Raonic et al., 2024), Complex neural operator (Tiwari et al., 2023), Wavelet neural operator (Gupta et al., 2021; Tripura and Chakraborty, 2023; Rani et al., 2023), Laplacian neural operator (Pang et al., 2020; Cao et al., 2024), RiemannONets (Peyvan et al., 2024), Geometry informed neural operator (GINO) (Li et al., 2024b), Diffeomorphism Neural Operator (Zhao et al., 2024), Spectral neural operator (Fanaskov and Oseledets, 2023; Rafiq et al., 2022), operator transformer (Hao et al., 2023; Li et al., 2022; Shih et al., 2025), Local neural operator (Li et al., 2024a), and Peridynamic neural operators (Jafarzadeh et al., 2024) are all trained using a vast amount of simulation data. The physics-informed variant of these methods also needs some amount of simulation data (perhaps in a coarse domain) for training. Moreover, as we show in this paper, the current leading architectures that can be trained without any simulation data are not efficient in practice for initial/boundary condition generalization.
@@
For learning generalized solutions of IBVPs, we develop the Physics-Informed Neural Transformer Operator (PINTO) that learns a map between the initial/boundary function space and the PDE’s solution function space. The primary novelty is introducing a cross-attention unit that modulates the internal representation of the domain’s space-time coordinates with information from the initial/boundary function, so that a generalized solution is obtained for any unseen initial/boundary function. Our model is trained with physics loss, without using simulation data.
@@
The applications that can benefit from the developed architecture are vast. For example, it could be used as surrogate models for faster inference (Chennault et al., 2021; Arcucci et al., 2020) and as solvers (Wu et al., 2021; Fablet et al., 2021) in data assimilation and inverse problems. It can replace traditional CFD solvers in applications that require fluid flow simulations such as autonomous path planning (Chowdhury et al., 2023), biomedical fluid dynamics (Kelley and Thomas, 2023), mask design (Mittal et al., 2023), and wake modeling in wind farms (Breton et al., 2017).
@@ 1.1 Related work:
Neural operators to learn maps between infinite-dimensional input and output spaces have gained tremendous popularity in recent years. Concurrent work on branch-trunk network DeepONet architecture, discretization-dependent convolutional neural operators, and discretization-invariant Fourier neural operators has led the way. In the DeepONet architecture, a trunk net encodes the coordinates of the space-time domain of the PDE, and a branch net encodes the discrete input function space (Lu et al., 2021a). The representations are merged using a Hadamard product. This operator was trained using simulation data. Later work extended the setup to situations where data availability is sparse by introducing PDE-residual loss as a physics-informed regularizer during DeepONet training (Wang et al., 2021, 2022b; Goswami et al., 2023). Extensions have been reported to multiple input functions (Jin et al., 2022), efficient numerics (Mandl et al., 2024), and training using the variational form of PDE residuals (Goswami et al., 2022).
@@
The Fourier neural operator (FNO) is a discretization invariant neural operator that is emerging as a potential universal approximator (Li et al., 2020a; Kovachki et al., 2021a, 2023a). Using convolutions within the spectral domain facilitated by the Fast Fourier Transform, FNOs learn the mapping between function spaces. However, they require a substantial amount of simulation data for training purposes. Enhancements in FNOs have been made by incorporating PDE-residual loss in their training, although they still require simulation data, but with reduced resolution (Li et al., 2024c). Deep implicit FNO has been used to learn the mapping between loading conditions and the response of heterogeneous materials without predefined constitutive models or microstructure measurements (You et al., 2022b). An FNO-based architecture has been developed to simulate a highly complex multiphase $\text{CO}_{2}$ water problem with wide ranges of porosity heterogeneity, anisotropy, reservoir conditions, injection configurations, flow rates, and multiphase flow properties using large simulation datasets (Wen et al., 2022). Fourier neural operator has been used for full wave inversion to infer subsurface structure information from seismic waveform data (Zhu et al., 2023).
@@
Transformers have been suggested as neural operators because they are inherently capable of mapping functional spaces over irregular and nonuniform grids. Data-driven operator learning has been conducted using an attention mechanism reminiscent of Galerkin methods (Cao, 2021). With this Galerkin-like attention, an encoder-decoder transformer model was developed to handle both uniform and varied discretization grids (Li et al., 2022). To solve complex PDEs involving multiple input functions and irregular grids, a general neural operator transformer Hao et al. (2023) was introduced. The authors implemented a heterogeneous normalized attention mechanism to decrease the computational cost of calculating attention scores and introduced a geometric gating mechanism to address multiscale problems.
Recently, an approach was proposed to solve PDEs for multiple initial and boundary conditions without retraining the model using modifications to the gradient descent algorithm (Navaneeth et al., 2024). They encode the prior knowledge of the governing equations through a stochastic projection-based gradient (Navaneeth and Chakraborty, 2023) that is computationally efficient in evaluating gradients for high-dimensional problems. Still, it comes with the cost of approximation and randomness in evaluating gradients. Their architecture learns the map between the initial conditions and the solution space.
@@
Though DeepONet achieved a wide range of applicability in approximating nonlinear operators, it has issues with scalability for high-dimensional PDEs due to data requirements. Also, the discrete input function requires data on a pre-defined grid that makes generalization difficult. FNOs and other recent neural operators mentioned in the introduction cannot be trained reliably without simulation data, restricting their applications to users without access to simulation data. The attention mechanism is a promising route for developing neural operators, and we utilize the same in our work. As a reference, we used the solution obtained by a physics-informed DeepONet (PI-DeepONet) to benchmark the performance of our proposed model.
@@ 1.2 Key contributions
Our main contributions are: (i) development of a new cross-attention mechanism, grounded in neural operator theory, for an efficient neural operator that generalizes the solution of initial boundary value problems for unseen initial/boundary conditions; (ii) training of the neural operator is using only physics loss, without data from numerical simulations; (iii) better simulation of solution for challenging benchmark test cases including the advection equation, Burger’s equation, Navier Stokes Equations in three different test cases; and (iv) ability to simulate solutions outside the temporal domain used for training. The success of our new neural operator is attributable to the transformation of the PDE solution’s domain query point (spatio-temporal coordinates) into a boundary-aware vector through multiple cross-attention units (iterative kernel integral operator).
In the remaining sections, we first describe the mathematical problem statement, neural operator theory, and practical implementation of our new architecture (Sect. 2). Next, we apply the new network for the $1$D Advection, $1$D nonlinear Burgers, the Kovasznay flow, the Beltrami flow, and the lid-driven cavity flow test cases (Sect. 3). We compare our results with other leading physics-informed neural operator architectures that can be trained without simulation data (i.e., PI-DeepONet) and showcase the advantages of the new model for initial and boundary condition generalization. In Sect. 4, we discuss the computational complexity, difference of PINTO from other physics-informed neural networks and neural operators, how to leverage our new cross-attention unit in other operators and limitations of PINN models. Finally, we conclude with a summary and directions for future work (Sect. 5).
@@ 2.1 Neural Operator Definition and Loss Function
where $h\in\mathcal{H}\subseteq\mathbb{R}^{s}$ is the $s$-dimensional solution field from the solution space $\mathcal{H}\subset L^{2}(\Omega,\mathcal{R}^{s})$, $\mathcal{N}$ is a general nonlinear differential operator with spatial and temporal partial derivatives, $X\in\Omega\subseteq\mathbb{R}^{d}$ is a $d$-dimensional coordinate from the spatiotemporal domain $\Omega$, $\mathbf{\alpha}$ is the PDE’s parameter vector, $f$ is a forcing term, $\mathcal{B}$ is the initial/boundary operator, $X_{b}\in\partial\Omega\subseteq\mathcal{R}^{d}$ is the $d$-dimensional initial/boundary coordinate from the domain’s boundary $\partial\Omega$ and $b$ is the imposed initial/boundary condition vector.
The initial boundary value problem can be expressed as follows. Let $\mathcal{A}\subset L^{2}(\partial\Omega,\mathcal{R}^{d})$ be the functional space of the initial/boundary condition. For an imposed initial/boundary condition $b\in\mathcal{A}$, there is a solution $h\in\mathcal{H}$ that satisfies the PDE (1). Thus, there exists a map $\mathcal{G}:\mathcal{A}\rightarrow\mathcal{H}$ that for every observed pair $(b,h)$ satisfies $h=\mathcal{G}(b)$. The solution field is accessible at any $X\in\Omega$ as $h(X)=\mathcal{G}(b)(X)$. We develop a parametrized neural operator $\mathcal{G}_{\theta}(X,b;\Theta^{*}),\Theta^{*}\in\mathcal{R}^{p}$ that approximates $\mathcal{G}$ and generalizes to predict the correct $h(X)$ for any $b\in\mathcal{A}$. Here, $p$ is the dimension of the parameter vector of the neural network.
@@
$\mathcal{B}(\mathcal{G}_{\theta}(X,b;\Theta^{*}),X_{b})=b\text{ on }\partial\Omega\,\forall\,b\in\mathcal{B}.$ (2b)
@@
$\min_{\Theta}\,\sum_{k=1}^{K}\{\frac{\lambda_{1}}{N_{c}}\sum_{j=1}^{N_{c}}|f_{c,\,k}^{j}-\mathcal{N}(\mathcal{G}(\Theta;X_{c,\,k}^{j});\alpha)|^{2}+\frac{\lambda_{2}}{N_{ib}}\sum_{j=1}^{N_{ib}}|b_{ib,\,k}^{j}-\mathcal{B}(\mathcal{G}(\Theta;X_{ib,\,k}^{j}))|^{2}\}\,,$ (3)

### v3

# PINTO: Physics-informed transformer neural operator for learning generalized solutions of partial differential equations for any initial and boundary condition

@@ Abstract
Applications in physics, engineering, mechanics, and fluid dynamics necessitate solving nonlinear partial differential equations (PDEs) with different initial and boundary conditions. Operator learning, an emerging field, solves these PDEs by employing neural networks to map the infinite-dimensional input and output function spaces. These neural operators are trained using data (observations or simulations) and PDE residuals (physics loss). A key limitation of current neural methods is the need to retrain for new initial/boundary conditions and the substantial simulation data required for training. We introduce a physics-informed transformer neural operator (named PINTO) that generalizes efficiently to new conditions, trained solely with physics loss in a simulation-free setting. Our core innovation is the development of iterative kernel integral operator units that use cross-attention to transform domain points of PDE solutions into initial/boundary condition-aware representation vectors, supporting efficient and generalizable learning. PINTO is utilized to simulate important equations used in fluid mechanics, physics and engineering applications: advection, Burgers, and steady and unsteady Navier-Stokes equations (three flow scenarios). We show that under challenging unseen conditions, the relative errors compared to analytical or numerical (finite difference and volume) solutions are low, merely 20% to 33% of those obtained by other leading physics-informed neural operator methods. Furthermore, PINTO accurately solves advection and Burgers equations at time steps not present in the training points, an ability absent for other neural operators. The code is accessible at https://github.com/quest-lab-iisc/PINTO.
@@
Nonlinear mapping , Operator learning , Transformer operator , Scientific machine learning , Physics-informed deep learning , Initial and boundary generalization , AI for PDEs
journal: Computer Physics Communications
{highlights}
New iterative integral kernel neural operator for learning generalized PDE solutions
Cross attention units transform solution domain representation to be boundary aware
Lifting and projection layers to map function spaces for efficient representation
Five challenging test cases demonstrate utility and improvement over leading models
Low relative error and ability to find solutions outside training examples
@@ 1 Introduction
Solving partial differential equations (PDEs) with different initial and boundary conditions is essential to understand and characterize the dynamics of natural and engineering systems. Traditionally, numerical methods (e.g., finite difference, finite element, finite volume, or spectral) are employed to solve PDEs. In recent decades, scientists have created high-performance computing software designed to solve PDEs, enabling their use as forward models in engineering design [1, 2, 3, 4]. However, for a new set of initial and/or boundary conditions, the numerical method must be applied afresh to solve the PDE as a new problem by spending significant compute resources. In contrast, contemporary artificial intelligence tasks in computer vision and natural language processing involve running a trained model (inference), which is computationally less expensive, for new scenarios. Those fields have benefitted from neural models that generalize to new situations well. The promise and application scope of such PDE solving models that could be trained once and used for fast simulation in new situations is appealing. Thus, over the past several years, the development of neural models for simulating systems governed by PDEs has picked up pace. Our goal is to develop a neural model for learning generalized solutions of PDEs for any initial and boundary condition.
Deep neural networks to solve PDEs are broadly classified into two categories: (i) neural operators that learn the mapping between function spaces (e.g., input function of say parameters and output function of the solution) such as DeepONets, Fourier neural operators (FNO), physics-informed neural operators (PINO), and graph neural operators (GNO) [5, 6, 7, 8, 9] and (ii) neural networks that learn a map between the PDE’s space-time domain and the solution function [10, 11, 12, 13, 14]. This distinction underscores the primary character of the input and output spaces used in these approaches. The latter is further classified into three (i) Physics-Informed Neural Networks (PINNs), (ii) Physics-Guided Neural Networks (PGNN) and (iii) Physics-Encoded Neural Networks (PENNs) [15].
All approaches can be trained using a combination of simulation data and PDE-residuals. The PGNN class of neural networks is trained primarily using simulations. They require a large data set of simulated (e.g., by solving PDEs numerically using traditional methods) or observed (experimental data) input and output. The PINN class of neural networks imposes the underlying physics by modifying the training objective to include the residuals of the PDEs. The residuals are calculated through automatic differentiation or through the discrete form of the PDEs in either weak or strong form [16, 17]. Energy-based loss functions have also been proposed [18, 19, 20]. The PENN class of neural networks attempts to impose known physics as part of the architecture itself, such as making the activation function learnable.
@@
PGNN, PINN and PENN learn the map for one instance of the PDE and require re-training for solving with new initial/boundary conditions, much like traditional numerical solvers. On the other hand, neural operators can learn a map between an initial/boundary function and a solution function, thereby achieving generalization to unseen conditions for a particular PDE. However, many of the current leading neural operators such as the Fourier neural operator (FNO) [6, 21, 22, 23, 24, 25, 26, 27, 28, 29], DeepONet [30, 7, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], Graph neural operator [5, 42], Convolutional neural operators [43], Complex neural operator [44], Wavelet neural operator [45, 46, 47], Laplacian neural operator [48, 49], RiemannONets [50], Geometry informed neural operator (GINO) [51], Diffeomorphism Neural Operator [52], Spectral neural operator [53, 54], operator transformer [55, 56, 57], Local neural operator [58], and Peridynamic neural operators [59] are all trained using a vast amount of simulation data. They resemble the PGNN framework; however, due to the fundamental difference in the input-output spaces, neural operators are generalizable.
Recently, physics-informed variants of neural operators (effectively combining the two major categories of neural models to solve PDEs) have been proposed [16, 60, 61]. These methods also need some amount of simulation data (perhaps in a coarse domain) for training. Moreover, as we show in this paper, the current leading architectures that can be trained without any simulation data are not efficient in practice for initial/boundary condition generalization.
@@
For learning generalized solutions of PDEs, we develop the Physics-Informed Neural Transformer Operator (PINTO) that learns a map between the initial/boundary function space and the PDE’s solution function space. The primary novelty is introducing a cross-attention unit that modulates the internal representation of the domain’s space-time coordinates with information from the initial/boundary function, so that a generalized solution is obtained for any unseen initial/boundary function. Our model is trained with physics loss computed by automatic differentiation, without using simulation data.
@@
The applications that can benefit from the developed architecture are vast. For example, it could be used as surrogate models for faster inference [62, 63] and as solvers [64, 65] in data assimilation and inverse problems. It can replace traditional CFD solvers in applications that require fluid flow simulations such as autonomous path planning [66], biomedical fluid dynamics [67], mask design [68], and wake modeling in wind farms [69, 70].
@@ 1.1 Related work:
Neural operators to learn maps between infinite-dimensional input and output spaces have gained tremendous popularity in recent years. Concurrent work on branch-trunk network DeepONet architecture, discretization-dependent convolutional neural operators, and discretization-invariant Fourier neural operators has led the way. In the DeepONet architecture, a trunk net encodes the coordinates of the space-time domain of the PDE, and a branch net encodes the discrete input function space [7]. The representations are merged using a Hadamard product. This operator was trained using simulation data. Later work extended the setup to situations where data availability is sparse by introducing PDE-residual loss as a physics-informed regularizer during DeepONet training [30, 71, 36]. Extensions have been reported to multiple input functions [72], efficient numerics [73], and training using the variational form of PDE residuals [35].
@@
The Fourier neural operator (FNO) is a discretization invariant neural operator that is emerging as a potential universal approximator [6, 27, 9]. Using convolutions within the spectral domain facilitated by the Fast Fourier Transform, FNOs learn the mapping between function spaces. However, they require a substantial amount of simulation data for training purposes. Enhancements in FNOs have been made by incorporating PDE-residual loss in their training, although they still require simulation data, but with reduced resolution [74]. Deep implicit FNO has been used to learn the mapping between loading conditions and the response of heterogeneous materials without predefined constitutive models or microstructure measurements [24]. An FNO-based architecture has been developed to simulate a highly complex multiphase $\text{CO}_{2}$ water problem with wide ranges of porosity heterogeneity, anisotropy, reservoir conditions, injection configurations, flow rates, and multiphase flow properties using large simulation datasets [22]. Fourier neural operator has been used for full wave inversion to infer subsurface structure information from seismic waveform data [75].
@@
Transformers have been suggested as neural operators because they are inherently capable of mapping functional spaces over irregular and nonuniform grids. Data-driven operator learning using an attention mechanism reminiscent of Galerkin methods has been proposed [76]. With this Galerkin-like attention, an encoder-decoder transformer model was developed to handle both uniform and varied discretization grids [56]. To solve complex PDEs involving multiple input functions and irregular grids, a general neural operator transformer [55] was introduced. The authors implemented a heterogeneous normalized attention mechanism to decrease the computational cost of calculating attention scores and introduced a geometric gating mechanism to address multiscale problems.
Recently, several approaches have been proposed to solve PDEs without using simultation data. For example, a finite element-based physics-informed operator learning framework has been developed with a Galerkin discretized weak formulation and an implicit Euler time integration scheme for temporal discretization as a loss function to incorporate physics into the network [16]. Similarly, the Lippmann-Schwinger operator in Fourier space has been used to construct physical constraints, effectively eliminating Automatic Differentiation [60]. The prior knowledge of the governing equations has been encoded through a stochastic projection-based gradient [77] to solve PDEs for multiple initial and boundary conditions without retraining the model. Still, these approaches come with the cost of approximation and randomness in evaluating gradients.
@@
Though DeepONet achieved a wide range of applicability in approximating nonlinear operators, it has issues with scalability for high-dimensional PDEs due to data requirements. Also, the discrete input function requires data on a pre-defined grid that makes generalization difficult. FNOs and other recent neural operators (see Section 1) cannot be trained reliably without simulation data, restricting their applications to users without access to simulation data. The attention mechanism is a promising route for developing neural operators, and we utilize the same in our work. As a reference, we used the solution obtained by a physics-informed DeepONet (PI-DeepONet) to benchmark the performance of our proposed model.
@@ 1.2 Key contributions
Our main contributions are: (i) development of a new cross-attention mechanism, grounded in neural operator theory, for an efficient neural operator that generalizes the solution of initial boundary value problems for unseen initial/boundary conditions; (ii) training of the neural operator uses only physics loss, without data from numerical simulations; (iii) better simulation than leading physics informed neural operator models for challenging benchmark test cases including the advection equation, Burger’s equation, and Navier Stokes Equations in three different test cases; and (iv) ability to simulate solutions outside the temporal domain used for training. The success of our new neural operator is attributable to the transformation of the PDE solution’s domain query point (spatio-temporal coordinates) into a boundary-aware vector through multiple cross-attention units (iterative kernel integral operator).
In the remaining sections, we first describe the mathematical problem statement, the neural operator theory, and the practical implementation of our new architecture (Sect. 2). Next, we apply the new network for the $1$D Advection, $1$D nonlinear Burgers, the Kovasznay flow, the Beltrami flow, and the lid-driven cavity flow test cases (Sect. 3). We compare our results with other leading physics-informed neural operator architectures that can be trained without simulation data (i.e., PI-DeepONet) and showcase the advantages of the new model for initial and boundary condition generalization. In Sect. 4, we discuss the computational complexity, differnce of PINTO from other physics-informed neural networks and neural operators, how to leverage our new cross-attention unit in other operators and limitations of PINN models. Finally, we conclude with a summary and directions for future work (Sect. 5).
Table 1: List of symbols with their meaning used in the present paper.
$\mathcal{N}$ Nonlinear differential operator
$\mathcal{B}$ Initial/boundary operator
$\mathcal{H}$ Solution space
$\Omega,\,\partial\Omega$ Spatio-temporal domain, and domain’s boundary
$X,\,X_{b}$ $d-$dimensional spatio-temporal coordinate in $\Omega$ and $\partial\Omega$ respectively
$\mathcal{A}$ Functional space of initial and boundary conditions
$h$ $s-$dimensional solution field in $\mathcal{H}$
$f,\,b$ Forcing term and Dirichlet boundary conditions respectively
$\alpha$ PDE’s parameter vector
$\mathcal{G}$ Map between initial and boundary conditions functional space $\left(\mathcal{A}\right)$ to solution space $\left(\mathcal{H}\right)$
$\mathcal{G}_{\theta}$ Neural operator that approximates $\mathcal{G}$ with $\theta$ as trainable parameters
$\Theta,\,\Theta^{*}$ Trainable parameters $\theta$ and its optimal value
$p_{n}$ Number of parameters in the neural network (number of elements in $\Theta$)
$N_{c}$ Number of collocation points
$N_{ib}$ Number of initial and boundary points
$\lambda_{i}$ Weights assigned to the $i$th term in the loss function
$\mathcal{Q}$ Projection operator that projects the learned representation from iterative kernel integration to solution space
$\mathcal{P}$ Lifting operator that lifts input to a higher dimensional space
$\mathcal{W}_{t}$ Local linear operator
$\mathcal{K}_{t}$ Nonlinear iterative kernel integral operator
$\gamma_{t}$ Bias term
$\sigma$ Nonlinear activation function
$\mu_{t}$ Hidden representation
$D_{t}$ Domain of hidden representation
$A_{h},\,B_{h},\text{ and }R_{h}$ Linear transformation in an attention head
$m$ Dimension of the encoded vector
$\zeta_{i}$ Attention score corresponds to $i^{th}$ value token
$L$ Length of discretized initial/boundary points
$x,y,t$ Space and time coordinates in the test cases.
$u,v,p$ $x,\,y$-directional velocities, and pressure in Navier Stokes test cases
$\beta$ Constant advection speed in advection test case
$\omega_{i},\,a_{i},\,\phi_{i},\,\Delta_{x}$ wave number, amplitude, phase angle, and domain size
in sinusoidal wave equation for initial conditions of advection test case
$\nu$ Kinematic viscosity in Navier Stokes test cases
@@ 2.1 Neural Operator Definition and Loss Function
where $h\in\mathcal{H}\subseteq\mathbb{R}^{s}$ is the $s$-dimensional solution field from the solution space $\mathcal{H}\subset L^{2}(\Omega,\mathbb{R}^{s})$, $\mathcal{N}$ is a general nonlinear differential operator with spatial and temporal partial derivatives, $X\in\Omega\subseteq\mathbb{R}^{d}$ is a $d$-dimensional coordinate from the spatiotemporal domain $\Omega$, $\mathbf{\alpha}$ is the PDE’s parameter vector, $f$ is a forcing term, $\mathcal{B}$ is the initial/boundary operator, $X_{b}\in\partial\Omega\subseteq\mathbb{R}^{d}$ is the $d$-dimensional initial/boundary coordinate from the domain’s boundary $\partial\Omega$ and $b$ is the imposed initial/boundary condition vector. Table 1 contains the list of symbols used in this paper with their definitions.
The initial boundary value problem can be expressed as follows. Let $\mathcal{A}\subset L^{2}(\partial\Omega,\mathbb{R}^{d})$ be the functional space of the initial/boundary condition. For an imposed initial/boundary condition $b\in\mathcal{A}$, there is a solution $h\in\mathcal{H}$ that satisfies the PDE (1). Thus, there exists a map $\mathcal{G}:\mathcal{A}\rightarrow\mathcal{H}$ that for every observed pair $(b,h)$ satisfies $h=\mathcal{G}(b)$. The solution field is accessible at any $X\in\Omega$ as $h(X)=\mathcal{G}(b)(X)$. We develop a parametrized neural operator $\mathcal{G}_{\theta}(X,b;\Theta^{*}),\Theta^{*}\in\mathbb{R}^{p_{n}}$ that approximates $\mathcal{G}$ and generalizes to predict the correct $h(X)$ for any $b\in\mathcal{A}$. Here, ${p_{n}}$ is the dimension of the parameter vector of the neural network.
@@
$\mathcal{B}(\mathcal{G}_{\theta}(X,b;\Theta^{*}),X_{b})=b\text{ on }\partial\Omega\,\forall\,b\in\mathcal{A}.$ (2b)
@@
$\min_{\Theta}\,\sum_{k=1}^{K}\{\frac{\lambda_{1}}{N_{c}}\sum_{j=1}^{N_{c}}|f_{k}^{j}-\mathcal{N}(\mathcal{G}(\Theta;X_{k}^{j});\alpha)|^{2}+\frac{\lambda_{2}}{N_{ib}}\sum_{j=1}^{N_{ib}}|b_{k}^{j}-\mathcal{B}(\mathcal{G}(\Theta;X_{b,\,k}^{j}))|^{2}\}\,,$ (3)